OpenStudy (anonymous):

Find the quotient. 6x - 3 / x^2 ÷ 5 / x a. 6x - 3 / 5x2 b. 6x - 3 / 5x c. 5x / 6x - 3 d. 5x2 / 6x - 3 help please.

OpenStudy (quietus):

Please use what @jim_thompson5910 has taught you :)

OpenStudy (quietus):

it;s all the same

OpenStudy (anonymous):

yes i know but the last part is confusing me still >.<

jimthompson5910 (jim_thompson5910):

what do you have so far?

OpenStudy (anonymous):

you switch to mutiply, and switch the last fraction

OpenStudy (anonymous):

then you cancel the Xs, and replace with 1.

OpenStudy (anonymous):

then, im lost. :C

jimthompson5910 (jim_thompson5910):

you are correct in saying that you flip and multiply \[\Large \frac{6x-3}{x^2} \div \frac{5}{x}\] \[\Large \frac{6x-3}{x^2} \bullet \frac{x}{5}\]

jimthompson5910 (jim_thompson5910):

I think you're stuck on the cancelling still

OpenStudy (anonymous):

crud :C

jimthompson5910 (jim_thompson5910):